OpenStudy (anonymous):

derivative of f(x)=sin²x(2x+1)

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

I'm on it!

OpenStudy (gorv):

is x(2x+1) are combined in sin 0r only x is there???? @cronch

OpenStudy (kainui):

So this looks like a combination of the chain rule and power rule. Show me your best guess @cronch and I'll help you figure out where you're going wrong.

OpenStudy (kainui):

It'll be more instructive if you try your best cronch, then once I point out where your mistake is, you won't have to worry about making that type of mistake again. =D

OpenStudy (calculusfunctions):

@Kainui I think @gorv made a good point. Is the question\[f(x) =\sin ^{2}[x(2x +1)]\]OR\[f(x)=(2x +1)\sin ^{2}x\]They are two different functions. @cronch which one are you asked to differentiate?

OpenStudy (kainui):

Only time will tell.

OpenStudy (anonymous):

I think distributing would be the best thing to do first.

OpenStudy (anonymous):

Would make it simpler.

OpenStudy (calculusfunctions):

"Only time will tell"?? @Kainui you sound like a fortune cookie lol.

OpenStudy (kainui):

Could use a trig identity to make sin^2(whatever) into a simple 1/2-1/2cos(2whatever) which can be foiled and also "easier" I guess. Haha, I'm hungry don't talk about cookies right now.

OpenStudy (anonymous):

It's sort of confusing on what the question is as well, if it's the second one that calculusfunctions pointed out then all you have to do is use the product rule really.

OpenStudy (calculusfunctions):

Yes @iambatman depending on the intended function, both your suggestions are the most efficient methods.

OpenStudy (calculusfunctions):

Apply the distributive property to simplify and then use the chain rule to differentiate, in the first case. Product rule for the second case.